# RL for Reasoning by Adaptively Revealing Rationales

### v1

@@ Abstract
We propose that reinforcement learning (RL) from partial expert demonstrations is not merely a training heuristic, but a promising framework for solving complex sequence generation tasks. Supervised fine-tuning (SFT) relies on dense ground-truth labels, which become increasingly costly as sequence length grows. RL, on the other hand, struggles with sparse rewards and a combinatorially large output space. We address this by introducing adaptive backtracking (AdaBack), a per-sample curriculum learning algorithm that reveals only a partial prefix of the target output during training. The supervision length is adjusted dynamically for each sample based on the model’s past reward signal, allowing it to incrementally learn to complete reasoning chains by conditioning on correct partial solutions. We investigate this intermediate regime between SFT and RL and argue that per-sample curriculum learning is more than a trade-off between efficiency and generality—it can succeed in tasks with long sequences of latent dependencies where SFT and RL both fail to generalize. Using a synthetic task with latent parity constraints, we show that our adaptive curriculum over partial answers reliably solves problems that are otherwise intractable. On mathematical reasoning benchmarks (MATH, GSM8k), we find that curriculum learning enables models to solve problems that RL alone cannot, acquiring new reasoning capabilities through incremental exposure to partial solutions.
@@ 1 Introduction
The reasoning capabilities of Transformers [53] have been extensively studied across domains such as mathematics [45, 15, 27, 35], algorithmic reasoning [54], and code generation [13]. These studies indicate that the reasoning performance of such models is significantly enhanced by training and inference mechanisms involving explicit reasoning traces or rationales, commonly known as scratchpad [40] or chain-of-thought [55].
However, acquiring large-scale, high-quality reasoning traces for specialized domains, such as mathematics, poses substantial challenges. Reinforcement learning (RL)-inspired methods have emerged as a promising solution to this challenge. By utilizing (potentially noisy) reward functions or verifiers (e.g., checking final answers to math problems), language models can generate novel reasoning traces, effectively using the model as an RL policy. For instance, the REINFORCE algorithm [56] has been employed by methods like STaR [60], generating multiple solutions per problem and selectively fine-tuning on the correct ones. More advanced methods utilize algorithms like Proximal Policy Optimization (PPO) [47] and, more recently, Group Relative Policy Optimization (GRPO) [48].
@@
These limitations motivate our investigation of a third regime: the space between supervised fine-tuning (SFT) and RL. We study whether breaking down long reasoning chains into partial rationales—and adaptively conditioning the model on only a prefix of the target during training—can make exploration tractable, and ultimately expand the class of problems that sequence models can learn to solve. This leads to our central research question:
Can reinforcement learning, when guided by adaptive partial supervision, teach models genuinely new reasoning capabilities?
@@
Specifically, can it enable discovery of solutions that were previously exponentially unlikely under the model’s initial distribution?
@@
To illustrate this clearly, consider a reasoning task consisting of $n$ sequential steps, each of which must be performed correctly for the task to succeed (e.g., proving a math theorem). For simplicity, assume the model performs each step correctly with a probability of $p=\Theta_{n}(1)$. Naively attempting the entire task yields a success probability of $p^{n}$, meaning that positive reinforcement signals would be exponentially infrequent—on average, every $\Theta(p^{-n})$ iterations—making standard RL impractical. To mitigate this, we propose adaptive guidance: initially, we reveal all but the final step of the solution from the training dataset, using RL solely for this final step. Consequently, the model receives feedback with probability $p$ every iteration. As performance improves, we progressively conceal more steps of the solution, training the model to complete increasingly larger segments, maintaining frequent positive feedback. Eventually, the entire reasoning chain is learned step-by-step, effectively transforming a complex search space with success probability $p^{n}$ into $n$ simpler sub-searches, each with success probability $\Theta(p)$. While this example includes simplifying assumptions, we empirically demonstrate this phenomenon’s practicality on a synthetic task described in Sec. 2.2.
@@
The scenario described above is simplified and idealized. In real-world problems, reasoning steps may not be clearly distinguishable, and datasets typically contain tasks with varying difficulties and solution lengths, making uniform partial reveals inefficient. To address these practical challenges, we propose a per-sample adaptive algorithm that dynamically adjusts the revealed portion of each solution based on its inherent difficulty. Specifically, we leverage the GRPO framework [48], which naturally estimates task difficulty by generating multiple rollouts per question and averaging the rewards. We term this method adaptive backtracking (AdaBack), detailed further in Sec. 2.
We note that backtracking strategies have precedents in the RL literature—for instance, the backtracking method of Salimans and Chen [43]. In language modeling, R3 [57] introduces a curriculum-based approach that progressively trains the model to complete reasoning chains from earlier positions. However, R3 ultimately performs static data augmentation, requiring domain-specific delimiters to segment reasoning chains—a limitation that impedes scalability and general applicability to datasets lacking clear step delimiters, such as MATH [27]. In contrast, AdaBack does not rely on clearly separated reasoning steps, instead employing a per-sample adaptive strategy suited for general discrete sequential data. Further comparisons with related methods are discussed in Sec. 4.
@@
  • We propose AdaBack, a per-sample adaptive curriculum learning algorithm that dynamically adjusts the amount of supervision during RL training based on reward feedback. AdaBack enables each sample to progress at its own rate, eliminating the need for manual curriculum staging or handcrafted schedules.
  • We show that AdaBack bridges the gap between SFT and RL, enabling learning in regimes where both fail. In particular, AdaBack transitions from full supervision to full exploration in a data-driven way, using a single threshold to govern difficulty progression.
  • On a synthetic parity task with sparse rewards, we demonstrate a separation result: AdaBack reliably solves the task, while SFT, RL, and their combination all fail.
  • On standard mathematical reasoning benchmarks (MATH, GSM8k), we show that AdaBack improves performance over standard RL and SFT+RL pipelines. We also show that AdaBack applied to base models often matches the performance of SFT-initialized counterparts.
  • We introduce two more variants of GSM8k—Base-7, which uses an unfamiliar numerical format never seen by the model during pretraining, and Tensor-2, which concatenates problems to increase reasoning depth. AdaBack achieves strong performance on both, demonstrating robust generalization to symbolic shifts and longer-horizon reasoning.
  • To study the potential limitations of the space of viable solutions and reasoning capabilities, in line with Yue et al. [59], we evaluate models using pass@k. We find that AdaBack significantly improves pass@k—especially when applied without SFT—suggesting that it enables models to discover novel solutions rather than simply reweighting existing ones.
@@ 2 Adaptive Backtracking
For RL algorithms like GRPO [48] or RLOO [6] that perform multiple rollouts for each sample, we dynamically adjust the supervision level so that the average sample reward across the rollouts stays close to a desired amount (for example, 50% of the answers are correct for each math question). For other RL algorithms, the criteria to change the supervision level can depend on other estimation of the state’s value.
@@
We denote random variables using capital letters. Let an input sample be represented as a sequence of tokens $X=(X_{1},X_{2},\dots,X_{\ell})$, corresponding to a natural language prompt, such as a question or problem description. The desired output is a chain-of-thought (CoT) style response $Y=(Y_{1},Y_{2},\dots,Y_{m})$, where each $Y_{t}$ represents a token or step in the model’s reasoning process.
@@ 2.1 Method
At each training step, given an input $X^{(i)}=(X^{(i)}_{1},\dots,X^{(i)}_{\ell_{i}})$ and its corresponding target output $Y^{(i)}=(Y^{(i)}_{1},\dots,Y^{(i)}_{m_{i}})$, we reveal only the first $k$ tokens of $Y$, $Y_{1:k}$, where $k=\lfloor\rho^{(i)}\cdot m_{i}\rfloor$ and $\rho^{(i)}\in[0,1]$ is a sample-specific supervision portion. The model is trained to generate the suffix conditioned on the input and the revealed prefix $\hat{Y}^{(i)}_{k+1:m_{i}}\sim P_{\theta}(\cdot\mid X^{(i)},Y^{(i)}_{1:k})$.
Figure 1: AdaBack Update Rule. At epoch $t$, we sample a supervision ratio $\rho^{(i)}_{t}\sim U(\rho_{\min}^{(i)},\rho_{\max}^{(i)})$ and condition the model’s generation on the question and the corresponding partial answer (shown in black text, with unrevealed content grayed out). If the expected reward is below a threshold $\tau$, we increase supervision in the next epoch by sampling from the red interval $\rho_{t+1}^{(i)}\sim U(\rho_{t}^{(i)},\rho_{\max}^{(i)})$. Otherwise, if the reward exceeds $\tau$, we reduce supervision and sample from $U(0,\rho_{t}^{(i)})$, the green interval, to make the task harder.
@@
For each training sample $i$, we maintain an interval $[\rho_{\min}^{(i)},\rho_{\max}^{(i)}]$ initialized as $[0,1]$, from which we uniformly sample the supervision portion $\rho^{(i)}_{t}$ for sample $i$ at epoch $t$. After generating a set of predictions and receiving an average reward $r^{(i)}_{t}$ (or other estimations of the state’s value if not using GRPO), we update this interval based on a fixed reward threshold $\tau$:
If $\text{If }r^{(i)}_{t}<\tau:\quad$ $\rho_{\min}^{(i)}\leftarrow\rho^{(i)}_{t}$
If $\text{If }r^{(i)}_{t}\geq\tau:\quad$ $\rho_{\max}^{(i)}\leftarrow\rho^{(i)}_{t},\quad\rho_{\min}^{(i)}\leftarrow 0.0$
@@ 2.2 Chain-of-Parities: A Synthetic Environment for Studying Reasoning
This design enforces a latent, step-wise structure over the output. The $Y_{i}$ values act as a “scratchpad”—arbitrary values that influence $Z_{i}$ through an accumulating parity computation. In essence, each $Z_{i}$ encodes the parity of the prefix $(X_{1},\ldots,X_{i},Y_{1},\ldots,Y_{i})$, making the correctness of $Z_{i}$ dependent on the accuracy of $Z_{i-1}$ and prior CoT steps. This recursive dependency mirrors real-world CoT tasks, where early mistakes cascade through the solution. Although there are $2^{L}$ valid outputs per input due to unconstrained $Y_{i}$, generating one randomly has a probability of just $2^{-L}$—analogous to sparse-reward regimes common in reasoning tasks.
@@
  • SFT fails: For small training sets, supervised training alone cannot learn this task. Specifically, the task involves learning $n-1$ parity functions of degree three, for which the theoretical sample complexity is well-studied [4, 33]. For example, in the statistical query (SQ) model [31], weakly learning a degree-$k$ parity from $L$ input bits requires at least $\Omega(L^{k-1})$ samples. Consequently, regular fixed-size neural networks trained via stochastic gradient descent (SGD) cannot even weakly learn a degree-three parity from only $n=\tilde{O}(L)$ [^1] We use $\tilde{O}$ to hide the logarithmic factors. samples [4].
@@
This setup enables the model to explore and learn the final reasoning step. Specifically, for each partial sequence $X_{1},\ldots,X_{L},Y_{1},\ldots,Z_{L-1}$, the model produces two valid continuations: $Y_{L},Z_{L}$ and $Y_{L}^{\prime},Z_{L}^{\prime}$, where $x^{\prime}=x\oplus 1$ denotes the complement of $x$. This variation allows gradient-based learners to infer that $Y_{L}$ directly influences $Z_{L}$: flipping $Y_{L}$ changes $Z_{L}$ while all other inputs remain fixed. Consequently, learning $Z_{L}$ reduces to learning a degree-two parity function, as one input ($Y_{L}$) is already learned, which has lower sample complexity than learning the degree-three parity function. In particular, degree-two parities can be learned with $\tilde{\Omega}(L)$ samples [21, 33]. Once the model masters the final step, the curriculum is adjusted to reveal a shorter solution prefix (up to $Z_{L-2}$), requiring the model to generate $Y_{L-1},Z_{L-1},Y_{L},Z_{L}$. Since it has already learned to generate $Y_{L},Z_{L}$, the focus shifts to learning the $(L-1)$-th step. This process is repeated iteratively by gradually shortening the hint until the model learns the full sequence. This progression demonstrates that there exist sample-size regimes where standard SFT + RL fails, but AdaBack successfully learns the task.
We note that the argument above is simplified, as the actual learning complexity depends on the specific learning model used. We will discuss such subtleties in Appendix B. Here, we focus on the empirical evidence supporting this phenomenon. We consider the case of $L=16$ with a training set size of $n=1024$, using the LLaMA 3.2 1B model. The reward function assigns a value of $1$ to fully correct sequences and $0.1$ to sequences that are syntactically valid (i.e., correct number of bits).
@@
This experiment highlights how curriculum learning enables guided exploration via partially correct sequences, effectively expanding the class of learnable problems in structured sequence modeling. It also establishes a clear separation between SFT + standard RL and AdaBack: while the former struggles with sparse rewards, AdaBack offers a principled middle ground, avoiding both the limitations of full supervision and the exploration challenges of standard RL.
Figure 2: Training Dynamics. Left: Training and test rewards along with supervision ratios throughout training. With AdaBack, LLaMA 3.2 1B successfully learns the task in under 700 iterations. Right: Training and test rewards for SFT+RL (red) plateau at 0.1, indicating that only the output format—learned during supervised pretraining—has been retained. Test reward for R3 [57] is shown in purple; it reaches only 0.8 reward after more than 16,000 iterations. R3 segments training examples at all whitespace positions and applies RL uniformly over these fragments, resulting in inefficiency due to its non-adaptive strategy.

### v2

# arXiv:2506.18110v2[cs.LG] 02 Mar 2026

@@ Abstract
Learning in the combinatorially large output space of sequence generation problems is challenging as providing expert demonstrations scales poorly with sequence length, and RL struggles with sparse rewards. Between dense demonstrations in supervised training and no demonstrations in reinforcement learning lies an underexplored regime: partial supervision. We ask whether some classes of sequence learning problems become efficiently learnable by exploiting this gap. We address this by introducing adaptive backtracking (AdaBack), a per-sample curriculum learning algorithm that reveals a partial prefix of the target output. The supervision length is adjusted dynamically for each sample based on the model’s past reward signal, allowing it to incrementally learn to complete reasoning chains by conditioning on correct partial solutions. We investigate this intermediate regime between SFT and RL and argue that per-sample curriculum learning is more than a trade-off between efficiency and generality—it can succeed in tasks with long sequences of latent dependencies where SFT and RL both fail to generalize. Using a synthetic task with latent parity constraints, we show that AdaBack reliably solves problems that are otherwise intractable. On three mathematical reasoning benchmarks, DeepScaleR, MATH, and GSM8k, we find that AdaBack enables models to solve problems that RL alone cannot, acquiring new reasoning capabilities through incremental exposure to partial solutions.
@@ 1 Introduction
The reasoning capabilities of Transformers [Vaswani et al., 2017] have been extensively studied across domains such as mathematics [Saxton et al., 2019, Cobbe et al., 2021, Hendrycks et al., 2021, Lewkowycz et al., 2022], algorithmic reasoning [Veličković et al., 2022], and code generation [Chen et al., 2021]. These studies indicate that the reasoning performance of such models is significantly enhanced by training and inference mechanisms involving explicit reasoning traces or rationales, commonly known as scratchpad [Nye et al., 2021] or chain-of-thought [Wei et al., 2023].
However, acquiring large-scale, high-quality reasoning traces for specialized domains, such as mathematics, poses substantial challenges. Reinforcement learning (RL)-inspired methods have emerged as a promising solution to this challenge. By utilizing (potentially noisy) reward functions or verifiers (e.g., checking final answers to math problems), language models can generate novel reasoning traces, effectively using the model as an RL policy. For instance, the REINFORCE algorithm [Williams, 1992] has been employed by methods like STaR [Zelikman et al., 2022], generating multiple solutions per problem and selectively fine-tuning on the correct ones. More advanced methods utilize algorithms like Proximal Policy Optimization (PPO) [Schulman et al., 2017] and, more recently, Group Relative Policy Optimization (GRPO) [Shao et al., 2024].
@@
These limitations motivate our investigation of a third regime: the space between supervised fine-tuning (SFT) and RL. We study whether breaking down long reasoning chains into partial rationales—and adaptively conditioning the model on only a prefix of the target solution during training—can make exploration tractable, and ultimately expand the class of problems that sequence models can learn to solve. This leads to our central research question:
@@
Can reinforcement learning, when guided by adaptive partial supervision, teach models genuinely new reasoning capabilities? Specifically, for sequences learning tasks of variable lengths, can it enable discovery of solutions that were previously exponentially (in sequence length) unlikely under the model’s initial distribution?
@@
To illustrate this clearly, consider a reasoning task consisting of $n$ sequential steps, each of which must be performed correctly for the task to succeed (e.g., proving a math theorem). For simplicity, assume the model performs each step correctly with a constant probability $p$. Naively attempting the entire task yields a success probability of $p^{n}$, meaning that positive reinforcement signals would be exponentially infrequent—on average, every $p^{-n}$ iterations—making standard RL impractical. To mitigate this, we propose adaptive guidance: initially, we reveal all but the final step of the solution from the training dataset, using RL solely for this final step. Consequently, the model receives feedback with probability $p$ every iteration. As performance improves, we progressively conceal more steps of the solution, training the model to complete increasingly larger segments, maintaining frequent positive feedback. Eventually, the entire reasoning chain is learned step-by-step, effectively transforming a complex search space with success probability $p^{n}$ into $n$ simpler sub-searches, each with success probability $\Theta(p)$. While this example includes simplifying assumptions, we empirically demonstrate this phenomenon’s practicality on a parity task described in Section 2.2.
@@
The scenario described above is simplified and idealized. In real-world problems, reasoning steps may not be clearly distinguishable, and datasets typically contain tasks with varying difficulties and solution lengths, making uniform partial reveals inefficient. To address these practical challenges, we propose a per-sample adaptive algorithm that dynamically adjusts the revealed portion of each solution based on its inherent difficulty. Specifically, we leverage the GRPO framework [Shao et al., 2024], which naturally estimates task difficulty by generating multiple rollouts per question and averaging the rewards. We term this method adaptive backtracking (AdaBack), detailed further in Section 2.
Backtracking strategies have precedents in RL—for example, the method of Salimans and Chen [2018]. In language modeling, R3 [Xi et al., 2024] applies a curriculum that progressively trains models to complete reasoning chains from earlier positions. Building on these ideas, we ask: how can such slicing be adapted per sample, guided by model performance at each iteration, without relying on a heuristically designed global curriculum?
@@
  • We propose AdaBack, a per-sample adaptive curriculum learning algorithm that dynamically adjusts the amount of supervision during RL training based on reward feedback. AdaBack enables each sample to progress at its own rate, eliminating the need for manual curriculum staging or handcrafted schedules, transitioning from full supervision to full exploration in a data-driven way.
  • We show that AdaBack bridges the gap between SFT and RL, enabling learning in regimes where both fail. On a synthetic parity task with sparse rewards, we demonstrate a separation result: AdaBack reliably solves the task, while SFT, RL, and their combination all fail.
  • On standard mathematical reasoning benchmarks (DeepScaleR, MATH, and, GSM8k), we show that AdaBack improves performance over standard RL and SFT+RL pipelines. We also show that AdaBack applied to base models often matches the performance of SFT-initialized counterparts.
  • We introduce further two variants of GSM8k: Base-7, which uses an unfamiliar numerical format never seen by the model during pretraining, and Tensor-2, which concatenates problems to increase reasoning depth. AdaBack achieves strong performance on both, demonstrating robust generalization to symbolic shifts and longer-horizon reasoning.
@@ 2 Adaptive Backtracking
For RL algorithms like GRPO [Shao et al., 2024] or RLOO [Ahmadian et al., 2024] that perform multiple rollouts for each sample, we dynamically adjust the supervision level so that the average sample reward across the rollouts stays close to a desired amount (for example, 50% of the answers are correct for each math question). For other RL algorithms, the criteria to change the supervision level can depend on other estimation of the state’s value.
@@
We denote random variables using capital letters. Let an input sample be represented as a sequence of tokens $X=(X_{1},X_{2},\dots,X_{\ell})$, corresponding to a natural language prompt, such as a question or problem description. The desired output is a chain-of-thought (CoT) style response $Y=(Y_{1},Y_{2},\dots,Y_{m})$, where each $Y_{t}$ represents a token or step in the model’s reasoning process. We assume we have a dataset of such $(X,Y)$ pairs.
@@ 2.1 Method
At each training step, given an input $X^{(i)}=(X^{(i)}_{1},\dots,X^{(i)}_{\ell_{i}})$ and its corresponding target output $Y^{(i)}=(Y^{(i)}_{1},\dots,Y^{(i)}_{m_{i}})$, we reveal the first $k$ tokens of $Y$, $Y_{1:k}$, where $k=\lfloor\rho^{(i)}\cdot m_{i}\rfloor$ and $\rho^{(i)}\in[0,1]$ is a sample-specific supervision portion. The model is trained to continue generation conditioned on the input and the revealed prefix $\hat{Y}^{(i)}_{k+1:m^{\prime}_{i}}\sim P_{\theta}(\cdot\mid X^{(i)},Y^{(i)}_{1:k})$, where $m^{\prime}_{i}-k$ is the length of the generated continuation.
Figure 1: AdaBack Update Rule. At epoch $t$, we sample a supervision ratio $\rho^{(i)}_{t}\sim U(\rho_{\min}^{(i)},\rho_{\max}^{(i)})$ and condition the model’s generation on the question and the corresponding partial answer (shown in black text, with unrevealed content grayed out). If the average reward is below a threshold $\tau$, we increase supervision in the next epoch by sampling from the red interval $\rho_{t+1}^{(i)}\sim U(\rho_{t}^{(i)},\rho_{\max}^{(i)})$. Otherwise, if the reward exceeds $\tau$, we reduce supervision and sample from $U(0,\rho_{t}^{(i)})$, the green interval, to make the task harder.
@@
For each training sample $i$, we maintain an interval $\rho_{\min}^{(i)},\rho_{\max}^{(i)}]$ initialized as $[0,1]$, from which we uniformly sample the supervision portion $\rho^{(i)}_{t}$ for sample $i$ at epoch $t$. After generating a set of predictions and receiving an average reward $r^{(i)}_{t}$ (or other estimations of the state’s value if not using GRPO), we update this interval based on a fixed reward threshold $\tau$:
$\text{If }r^{(i)}_{t}<\tau:\quad$ $\rho_{\min}^{(i)}\leftarrow\rho^{(i)}_{t}$
$\text{If }r^{(i)}_{t}\geq\tau:\quad$ $\rho_{\max}^{(i)}\leftarrow\rho^{(i)}_{t},\quad\rho_{\min}^{(i)}\leftarrow 0.0$
@@ 2.2 Chain-of-Parities: A Synthetic Environment for Studying Reasoning
This design enforces a latent, step-wise structure over the output. The $Y_{i}$ values act as a “scratchpad”—arbitrary values that influence $Z_{i}$ through an accumulating parity computation. In essence, each $Z_{i}$ encodes the parity of the prefix $(X_{1},\ldots,X_{i},Y_{1},\ldots,Y_{i})$, making the correctness of $Z_{i}$ dependent on the accuracy of $Z_{i-1}$ and prior chain-of-thought (CoT) steps. This recursive dependency mirrors real-world CoT tasks, where early mistakes cascade through the solution. Although there are $2^{L}$ valid outputs per input due to unconstrained $Y_{i}$, generating one randomly has a probability of just $2^{-L}$—analogous to sparse-reward regimes common in reasoning tasks.
@@
  • SFT fails: For small training sets, supervised training alone cannot learn this task. Specifically, the task involves learning n1n-1 parity functions of degree three, for which the theoretical sample complexity is well-studied [Abbe et al., 2023a, Kou et al., 2024]. For example, in the statistical query (SQ) model [Kearns, 1998], weakly learning a degree-$k$ parity from $L$ input bits requires at least $\Omega(L^{k-1})$ samples. Consequently, regular fixed-size neural networks trained via stochastic gradient descent (SGD) cannot even weakly learn a degree-three parity from only $n=\tilde{O}(L)$ [^1] We use $\tilde{O}$ to hide the logarithmic factors. samples [Abbe et al., 2023a].
@@
This setup enables the model to explore and learn the final reasoning step. Specifically, for each partial sequence $X_{1},\ldots,X_{L},Y_{1},\ldots,Z_{L-1}$, the model produces two valid continuations: $Y_{L},Z_{L}$ and $Y_{L}^{\prime},Z_{L}^{\prime}$, where $x^{\prime}=x\oplus 1$ denotes the complement of $x$. This variation allows gradient-based learners to infer that $Y_{L}$ directly influences $Z_{L}$: flipping $Y_{L}$ changes $Z_{L}$ while all other inputs remain fixed. Consequently, learning $Z_{L}$ reduces to learning a degree-two parity function, as one input ($Y_{L}$) is already learned, which has lower sample complexity than learning the degree-three parity function. In particular, degree-two parities can be learned with $\tilde{\Omega}(L)$ samples [Glasgow, 2024, Kou et al., 2024]. Once the model masters the final step, the curriculum is adjusted to reveal a shorter solution prefix (up to $Z_{L-2}$), requiring the model to generate $Y_{L-1},Z_{L-1},Y_{L},Z_{L}$. Since it has already learned to generate $Y_{L},Z_{L}$, the focus shifts to learning the (L1)(L-1)-th step. This process is repeated iteratively by gradually shortening the hint until the model learns the full sequence. This progression demonstrates that there exist sample-size regimes where standard SFT + RL fails, but AdaBack successfully learns the task.
We note that the argument above is simplified, as the actual learning complexity depends on the specific learning model used. We will discuss such subtleties in Appendix B. Here, we focus on the empirical evidence supporting this phenomenon. We consider the case of $L=16$ with a training set size of $n=1024$, using the Llama 3.2 1B model [Grattafiori et al., 2024]. The reward function assigns a value of $1$ to fully correct sequences and $0.1$ to sequences that are syntactically valid (i.e., correct number of bits).
@@
This experiment highlights how AdaBack enables guided exploration via partially correct sequences, effectively expanding the class of learnable problems in structured sequence modeling. It also establishes a clear separation between SFT + standard RL and AdaBack: while the former struggles with sparse rewards, AdaBack offers a principled middle ground, avoiding both the limitations of full supervision and the exploration challenges of standard RL.
Figure 2: Training Dynamics. Left: Training and test rewards along with supervision ratios throughout training. With AdaBack, Llama 3.2 1B successfully learns the task in under 700 iterations. Right: Training and test rewards for SFT+RL (red) plateau at 0.1, indicating that only the output format—learned during supervised pretraining—has been retained. Test reward for R3 [Xi et al., 2024] is shown in purple; it reaches only 0.8 reward after more than 16,000 iterations. R3 segments training examples at all whitespace positions and applies RL uniformly over these fragments, resulting in inefficiency due to its non-adaptive strategy.